### v1

# Review_48_hours.bib

Abstract
We collect papers forecasting wind turbine failures at least two days in advance. We examine the prediction time, methods, failed components, and dataset size. We investigate the effect of using standard SCADA data and data from additional sensors, such as those measuring vibration. We observe a positive correlation between dataset size and prediction time. In the considered cases, one may roughly expect a forecast for at least two days using a dataset of one turbine year and a forecast for two hundred days from a dataset of a hundred turbine years.
Keywords: long range forecast, degradation time, remaining useful lifetime.
1   Introduction
This study aims at systematizing approaches that predict failures in wind turbines at least two days in advance. This requirement is common among wind turbine operators, see, e.g., the wind energy challenge raised by Energias de Portugal [EDP2018]. This is in line with the fact that powerful wind turbines are big and complicated structures with high repair and downtime costs, and require skilled personnel and special equipment for maintenance. Therefore, it takes some time to decide which repair team is needed and which spare parts need to be found and delivered. Moreover, the minimal required prediction time may further increase because there is a trend to have bigger and more powerful wind turbines [wiser2021expert].
Apart from the prediction time, the size of the available dataset may also influence the outcome. Some methods work better for small datasets, while others work for large datasets, and a minimal size of the dataset required for analysis may exist. Unfortunately, it is difficult to obtain and curate large datasets with a consistent and good quality [Kusiak2016]. Therefore, it is interesting to estimate to what extent the dataset size influences the prediction time of fault detection methods.
We propose using turbine years (TY) to measure the size of a dataset. One obtains TY by multiplying the number of working turbines by the number of years they worked. The term TY is similar to labor hour, machine hour, and horsepower, and its purpose is the same – standardization and the possibility to compare. The term TY exists and is used, for example, to compare the size of datasets [dao2019wind] or to quantify the harm of wind turbines to birds [mcclure2021eagle, ferrer2022significant], but we see a margin for increasing the popularity of this term.
The paper is organized as follows: Section 2 describes the methods we used to find relevant papers. Section 3 shows our results and findings, while Section 4 concludes the paper.
2   Methodology
The number of research articles on wind turbines is enormous and growing rapidly. A search for ”wind turbine” in Scopus in July 2024 gave us more than 100,000 indexed papers. Moreover, this number grows with acceleration. There were 5,000 papers indexed only in one year of 2013, while in 2023 this number of indexed papers was about 9,000. Limiting the search results only to the words ”wind” and ”turbine” found one after another with zero words in between (wind w/0 turbine), only to articles published in journals (no books and no proceedings), and only in the English language reduces the mentioned numbers to about 47,000, 1,600 and 4,800 correspondingly, with the most important factor being the proceedings exclusion. This abundance means that it is practically impossible even to read every article published within five recent years, and one must rely on some filtering. Therefore, we set our aim not to cover every paper but rather show popular solutions.
Our approach is twofold. On the one hand, we collected the papers from the literature that were relevant for a research project on fault prediction of wind turbines in the context of a specific case study with an open dataset, see Ref. [begun2024cost]. These papers we found using Google Scholar. On the other hand, we used Scopus to make a thorough ”scan” of articles published within the last five years that we found for a search phrase ”wind w/0 turbine and SCADA and signal”, where SCADA means Supervisory Control and Data Acquisition system, which is the name for standard installed sensors, and signal should be replaced with a term meaning a specific signal, for example, vibration. The use of Google Scholar is motivated by the fact that it is free of charge, counts citations well, thus showing approaches that are relevant to others, and is widely available outside academic institutions. The use of Scopus is motivated by the fact that it is a curated proprietary database that allows complicated searches, counts citations as well, and sorts by year even better.
The choice of SCADA and a signal was to limit the number of papers for deep reading and to check a popular hypothesis that modern techniques allow making good predictions using only SCADA data without dedicated and typically expensive additional sensors. The signal measures that we used in the Scopus search include the mentioned ”vibration” and also ”strain”, ”torque”, ”LiDAR”, and ”acoustic”. We also tried to search other signals, see Ref. [badihi_comprehensive_2022] for a comprehensive list of possibilities. However, they give either too rare results in combination with SCADA, e.g. ”thermography”, or too many false positives as ”electric”.
The most common is SCADA data with a 10-minute acquisition frequency. In principle, 5-, 1-minute, and other SCADA frequencies exist, but they are rare. The change from 10- to 1-minute frequency increases the amount of data 10 times but still does not allow us to monitor fast processes. A radically different amount of data appears if one considers vibrations because even a 1-minute frequency is only 1/(60 s) = 0.017 Hz, while vibration sensors measure values of the order of kHz. Therefore, we do not multiply the amount of more frequent SCADA data by a corresponding factor, but treat the SCADA data, for which vibration measurements are available, separately.
A turbine consists of many interrelated elements that form different groups. To classify failures into a small number of categories, we generalize compound terms based on our understanding. For example, we call a generator cooling failure discussed in Ref. [borchersen2016model] a generator failure because it addresses overheating of the generator, which has a different construction, sensors, and working modes than, e.g., a transformer, which also has a cooling. Another example is generator bearing and gearbox bearing failures discussed in Ref. [schlechtingen2011comparative], which we classify as bearing failures because these failures are rooted in bearings.
We analyze only the best predictions of the authors and disregard false alarms, failures that have not been found, absence of failure decision mechanism, validation or test, or both, and also accept the prediction if authors at least see a failure with their approach. The latter is particularly common for small datasets with TY<<<10. If several predictions are available, then we take an arithmetic mean. These are subjective choices. However, we provide citations to the discussed papers, and a reader can make his own classifications of the papers with long-term predictions that we found.
3   Results
The main result of this work is the list of found articles shown in tables 1 and 2 with a corresponding list of acronyms in Appendix, and figure 1, which shows the relation between the size of a dataset and prediction time. Tables are sorted according to the prediction time. Prediction days and the number of TY larger than one are rounded to integers.
Author Component Method TY Days
TY <<< 10, specific methods often used
Pei [pei2018wind] any anomaly FV, MR, kNN 5 150
Cui [cui2018anomaly] Gearbox NARX, WT, NN 6 105
Hu [hu2018prediction] Bearing WP, MA 0,8 92
Ferguson [ferguson2019standardisation] Gearbox Phys 2 45
Garan [garan_data-centric_2022] Bearing, Gearbox FI, HCF, MI, PCA, ICA, DT 10 30
Miele [Miele_2022] Gearbox, Generator, Bearings, Transformer GCA-MTS, LSTM-AE, CNN–LSTM, GMI, LRI 10 26
Bonacina [Bonacina_2022] Bearing PPS 10 18
Lázaro [lazaro2020determining] Gearbox GMC, HIn 3 14
Wang [wang2019early] Bearing RBM, MD, kMC 2 3
Chen [chen2018generalized] Bearing CombP, IF, SVM, RBFNN 5 2
Zhang [zhang2022dynamic] Gearbox, Converter, Rotor GRU, dAE, SVR, MD, dThr 0,4 2
Normal behavior model
Borchersen [borchersen2016model] Generator KF, CUSUM 129 300
Schlechtingen [schlechtingen2011comparative] Bearing, Stator AR, NN 21 119
Encalada-Dávila [encalada2021wind] Bearing NN, BayR 24 105
Tutivén [tutiven2022early] Bearing SVM, EWMA 107 90
Begun [begun2024cost] Hydraulic LinR, DI-CUSUM 10 44
Jankauskas [Jankauskas_2023] Gearbox GRU, LSTM 10 35
Rashid [rashid2020fault] Gearbox BagR 8 34
Barber [barber_enabling_2022] Hydraulic, Generator, Bearing LoMST, WHC-LOF, CUSUM 10 30
Yuan [yuan2019gearbox] Gearbox DBSCAN, RF, GBDT, XGBOOST, MD, QF 11 18
Kong [kong2020condition] Gearbox PPMC, CNN, GRU, EWMA 13 17
Wang [wang_operating_2022] Gearbox DBN, DDPG-SOM 10 16
Yan [yan2014condition] Bearing NN 3 16
Schlechtingen [schlechtingen2014wind] Hydraulic, Gearbox, Converter, Anemometer, Power ANFIS, FL 53 14
Liu [liu_condition_2022] Generator, Hydraulic, Transformer MR, LASSO, LARS 10 9
TY >>> 10, machine learning and basic statistical methods often used
McKinnon [mckinnon2021investigation] Pitch IF 30 450
Herp [herp2019assessment] Bearing WD, RNN, LSTM 36 245
Zaher [zaher2009online] Gearbox NN 52 180
Letzgus [letzgus2020change] Gearbox, Generator, Pitch, Hydraulic, Electrical KCPD 66 120
Zhao [zhao2016fault] Generator PCA, DBSCAN, SVM, ARIMA 47 44
Eriksson [eriksson2020machine] Gearbox, Generator, Bearing, Transformer, Hydraulic ”Kafka”: NN, PCA, LinR, DT, RF, L-SVC 10 25
Zhao [zhao2017fault] Generator PCA, SVM, SMOTE, DBSCAN, AOIn, SVM, NN, kNN, BayC 225 18
Chen [chen2013wind] Pitch ANFIS 65 14
Djeziri [djeziri2018hybrid] general approach PCA, BGraph, DegrV 3
Table 1: First author of a paper, failed component, method used for a solution, size of used dataset in TY, prediction in the number of days. The list of acronyms is in Appendix.
Author Component Method TY Days
Vibration sensors used
Rezamand [rezamand2020integrated] Bearing WT, WA 1360 328
Natili [natili2021multi] Bearing SVR, PCA 3 210
Carroll [carroll2019wind] Gearbox SVM, LogR, NN 300 165
Turnbull [turnbull2021combining] Gearbox, Generator SVM 2 98
Teng [turnbull2021combining] Bearing NN 0,6 26
Zhu [zhu2021condition] Gearbox DAE, PCA, SVM, kNN, LOF, IF 25 13
Cheng [cheng2018enhanced] Bearing PF, ANFIS 1 12
Xu [xu2021fault] Bearing dAE 11 10
Hu [hu2022health] Generator MA, QF 2 6
Table 2: The same as table 1 for the case of using additional vibration sensors.
Figure 1: Prediction time as a function of dataset size measured in turbine years. The dashed line is 2*TY. Labels denote four groups: predictions for small datasets with TY<<<10 (Specific), large datasets with TY>10absent10>10> 10 (ML/Statistic), normal behavior models (NBM), and datasets with both SCADA and vibration measures (Vibration).
We found 42 articles, that give forecasts for 2 days and more:
[…]
NBM is the group of methods that best follows the scaling. We assign a point to the group if authors build a model of normal behavior and then analyze the accumulated deviations from it. The group that does not follow the scaling is Vibration. The reason is that vibration gives access to fundamentally different information with much better time resolution than SCADA. Another reason is that we do not take into account the size of the vibration dataset, which is hard to estimate because authors often do not specify vibration frequencies and the time cuts they applied to reduce the amount of vibration data. The groups Specific and ML/Statistics are not strictly defined, and authors from these groups sometimes use the same or similar methods which may also overlap with NBM. However, we see a tendency to use advanced mathematical methods and sophisticated modeling for datasets with TY<<<10, while for larger datasets authors more often use the ”out of the box” machine learning methods for classification and have enough data to apply filters. The examples are the usage of Wiener process [hu2018prediction] and Gaussian mixture copula model [lazaro2020determining] for small datasets, while for large datasets typical methods include direct classification using isolation forest [mckinnon2021investigation] or a perceptron [zaher2009online] for differentiating between normal and abnormal states. Small datasets are also more often analyzed with special features designed by authors only for these datasets and not found in other works. These features are also called scores, factors or indicators, see the exact author’s namings in the list of Acronyms. One may notice that the list of acronyms is larger than the number of papers and only a few acronyms repeat. This means that a comparison of the performance of different methods is difficult, while there are not enough statistics for the methods.
It is hard to find more papers because of the variety of possible methods and their combinations. Authors also do not report their prediction time in a unified format that can be searched. The size of the dataset one has to calculate from the authors’ notice of the start and end of measures, which is also not unified or even not mentioned, and one has to calculate it from their figures. Therefore, one can not systematically search by these parameters.
Our systematic Scopus search for ”wind w/0 turbine and SCADA and vibration” brought us mainly the articles that we already have seen using Google Scholar, with the exception of Ref. [tutiven2022early]. The Scopus search for other sensors brought us only false positives. We found that the keywords strain and torque are used to examine the main structural elements such as foundation and tower of the turbine. These elements are designed for much longer services than other components of a turbine. The typical time scale is 20 years and there is also a large safety margin. Therefore, corresponding fatigue life evaluations can predict 20, as well as 200 years remaining life [mai2019prediction, zhao2023fatigue]. The ”LiDAR” as a searched signal gives articles with predictions no longer than 5 minutes, but clear clustering by amount of TY. Among 21 found articles, half analyze datasets with less than 1 TY and discuss mainly yaw and alignment, while the rest discuss mainly power reduction due to wake and aging. The ”acoustic” gives articles with only instant measures without prediction and for typically very small datasets starting from 0.026 TY. This particular number of TY originates from 4 turbines and 2 hours of measurements every night in February [iannace2019wind]. The predicted value is not the condition of the turbine, but the noise it makes.
Answering the question of whether SCADA and a dedicated sensor are always better than SCADA is cumbersome. First of all, there are no openly available datasets with SCADA and vibration data. Therefore, one has only a solution given by the analyzers of the dataset, which is not possible to check independently. The analyzers can be over-presenting positive results and under-publishing negative results. One can see from Fig. 1 how rare the points with TY>>>10, in other words, how difficult it is to get a large dataset, especially with vibration data. The owners of such datasets may be interested in proving that vibration sensors are worth the money already paid for them. The authors also did not ask this question themselves and were just showing the result of the analysis without checking whether one can get the same prediction with SCADA only. However, we found one paper, which shows that both SCADA and vibration data give comparable results [natili2021multi]. Our Figure 1 also suggests that having vibration data can be beneficial, but with enough data and some effort, one can get similarly good results using only SCADA data.
One can advocate the usage of rated power and production year with arguments similar to those for using TY. Turbines of a similar power should have a similar construction, hence similar failure modes and rates. There could also be a dependence on the calendar year when a dataset was analyzed. The rapid growth of computing power, the development of machine learning algorithms, and growing interest in condition monitoring may also be important factors. A fashion for using certain models may exist and change with time. Authors of Ref. [Montero_Jimenez_2020, Chatterjee_Dethlefs_2021] observe strong fluctuations and kinds of waves in the distribution of publications per model and in keywords. We do not have enough data to check this hypothesis. However, we see that the most common rated power we find in papers with our search methods is 2 MW, while a few papers discuss turbines with other rated power in the range from 600kW to 4 MW. The notable effect is that the keywords vibration, torque, and strain allow finding papers talking about structural elements like towers and foundations of large offshore wind turbines with the rated power of the order of 10 MW, see, e.g., Ref. [de2023long].
4   Conclusions
We found and classified papers that make long-term forecasts of wind turbine failures by prediction time, size of the dataset, failed component, and methods used for the solution. We see a positive correlation between the size of a dataset and prediction time with approximate linear scaling. With datasets of 1-100 TY, one may expect a 2-200 day forecast.
We do not observe any grouping by failed components in the found articles with long-term predictions. However, we see that most studies consider short-term predictions, with the exception of those focused on structural elements like towers and foundations, which are forecast for many years. We note also that vibration is measured for rotating elements like gearbox, generator, and bearing, which represent a significant amount but not all turbine components. The methods used for the forecasts are very versatile. However, we see a tendency to use specific methods and advanced modeling for datasets smaller than 10 TY, and more machine learning and statistical approaches with filtering for larger datasets.
We propose using TY to indicate the size of the used dataset for better search by this parameter and because of the development of different machine learning methods. Some methods require more data, while others work better on small datasets. Separation by TY allows looking at the models that work better for either of the cases, thus giving access to measuring performance and selection of the best suitable methods.
Appendix: List of Acronyms
ANFIS
adaptive neuro-fuzzy interference system
AIn
anomaly index
AOIn
anomaly operation index
AR
auto regression
AE
autoencoder
BagR
bagging regression
BayC
Bayesian classifier
BayR
Bayesian regularization
bi-LSTM
bidirectional LSTM
BGraph
Bond Graph
CombP
combination prediction
CNN
convolutional neural network
CUSUM
cumulative sum
DI-CUSUM
decision interval CUSUM
DT
decision tree
dAE
deep autoencoder
DBN
deep belief network
DDPG-SOM
deep deterministic policy gradient-self-organizing map
DAE
denoising auto-encoder
DBSCAN
density-based spatial clustering of applications with noise
DegrV
degradation velocity
EWMA
exponentially weighted moving average
XGBOOST
extreme gradient boosting
FI
feature importance
FV
feature vector
FL
fuzzy logic
GRU
gated recurrent unit
GMC
gaussian mixture copula model
GMI
global mahalanobis indicator
GBDT
gradient boosting DT
GCA-MTS
graph convolutional autoencoder for multivariate time series
HIn
health index
HCF
high correlation filter
ICA
independent component analysis
IE
information entropy
IF
isolation forest
kNN
k nearest neighbours
KF
Kalman filter
KCPD
kernel-based change-point detection
kMC
k-means clustering
LASSO
least absolute shrinkage and selection operator
LARS
least-angle regression
LinR
linear regression
L-SVC
linear supporct vector machine classifier
LoMST
local minimum spanning tree
LOF
local outlier factor
LRI
local residual indicator
LogR
logistic regression
LSTM
long short-term memory neural network
MD
Mahalanobis distance
MA
moving average
MLR
multivariate linear regression
MR
multivariate regression
MI
mutual information
NN
neural network that does not have a specific name
NARX
nonlinear autoregressive NN with exogenous inputs
PF
particle filtering
PPMC
Pearson prod-moment correlation
Phys
physics mased model
PPS
power predictive score
PCA
principal component analysis
QF
quantile filtering
RBFNN
radial basis function neural network
RF
random forest
RNN
recurrent neural network
RBM
restricted Boltzmann machine
SVM
support vector machine
SVR
support vector regression
SMOTE
synthetic minority over-sampling technique
WHC
Ward hierarchical clustering
WT
wavelet transform
WD
Weibull distribution
WA
weighted average
WP
Wiener process
[…]